# PolyLUT-Add: FPGA-based LUT Inference with Wide Inputs

### v1

@@ V Conclusion
We introduced PolyLUT-Add, a novel technique designed to enhance connectivity between neurons in LUT-based networks to deploy DNNs at the edge efficiently. By combining base PolyLUT models, our approach mitigates scalability issues associated with conventional implementations and significantly improves efficiency. Specifically, we demonstrated that by utilizing a configuration of $A=2$, PolyLUT-Add with a lower polynomial degree $D$ and fan-in $F$ are sufficient to achieve comparable accuracy to PolyLUT. Over our benchmarks, PolyLUT-Add achieved reductions in LUT consumption by factors of 1.3-7.7 with a 1.2-2.2 times decrease in latency. The PolyLUT-Add architecture enhances LUT-based neural network performance in terms of area efficiency and latency.
Future work could be developing a targeted optimization technique for the individualized adjustment of $F$, $A$ and $\beta$ parameters within each layer or neuron, aiming to substantially boost network accuracy, reduce latency, and optimize area efficiency.

### v2

@@ V Conclusion
We introduced PolyLUT-Add, a novel technique designed to enhance connectivity between neurons in LUT-based networks to efficiently deploy DNNs at the edge. By combining base PolyLUT models, our approach mitigates scalability issues associated with conventional implementations and significantly improves efficiency. Specifically, we demonstrated that by utilizing a configuration of $A=2$, PolyLUT-Add with a lower polynomial degree $D$ and fan-in $F$ are sufficient to achieve comparable accuracy to PolyLUT. Over our benchmarks, PolyLUT-Add reduced LUT consumption by factors of 2.0-13.9 with a 1.2-1.6 times decrease in latency. The PolyLUT-Add architecture enhances LUT-based neural network performance in terms of area efficiency and latency.
Future work could develop a targeted optimization technique for individually adjusting $F$, $A$, and $\beta$ parameters within each layer or neuron to substantially boost network accuracy, reduce latency, and optimize area efficiency.
VI Acknowledgments
Binglei Lou gratefully acknowledges financial support from Spatial Computational Learning.